OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, 7); Slope: - 2/3

OpenStudy (hoblos):

y-7=-2/3(x-1)

OpenStudy (anonymous):

that is not one of my answers??

OpenStudy (hoblos):

can you post the answers please

OpenStudy (anonymous):

y + 7 =-2/3(x-1) y -7 =- 2/3(x + 1) y + 7 =2/3(x + 1) y -7 =2/3(x-1)

OpenStudy (anonymous):

hobios looks right y-y1= m(x-x1)

OpenStudy (anonymous):

these are my answer to the problem: y + 7 =-2/3(x-1) y -7 =- 2/3(x + 1) y + 7 =2/3(x + 1) y -7 =2/3(x-1)

OpenStudy (anonymous):

are you sure it's this : Point: (1, 7); Slope: - 2/3

OpenStudy (hoblos):

mmm can you check the question again.. it seems there is a mistake in it

OpenStudy (anonymous):

yes those are the answers! :( i will have to check with teacher!

OpenStudy (anonymous):

simply ............y-7=-2/3(x-1) .now solve it .